### v1

Abstract
In recent years, climate extremes such as floods have created significant environmental and economic hazards for Australia, causing damage to the environment and economy and losses of human and animal lives. An efficient method of forecasting floods is crucial to limit this damage. Techniques for flood prediction are currently based on hydrological, and hydrodynamic (physically-based) numerical models. Machine learning methods that include deep learning offer certain advantages over conventional physically based approaches, including flexibility and accuracy. Deep learning methods have been promising for predicting small to medium-sized climate extreme events over a short time horizon; however, large flooding events present a critical challenge. We present an ensemble-based machine learning approach that addresses large-scale extreme flooding challenges using a switching mechanism motivated by extreme-value theory for long-short-term-memory (LSTM) deep learning models. We use a multivariate and multi-step time-series prediction approach to predict streamflow for multiple days ahead in the major catchments of Australia. The ensemble framework also employs static information to enrich the time-series information, allowing for regional modelling across catchments. Our results demonstrate enhanced prediction of streamflow extremes, with notable efficacy for large flooding scenarios in the selected Australian catchments. Through comparative analysis, our methodology underscores the potential for deep learning models to revolutionise flood forecasting across diverse regions.
keywords:
[…]
1 Introduction
In Australia, flooding is among the most devastating and costly natural disasters in terms of economic loss [1, 2], damage to natural habitats and environment, and loss of human and animal lives [3, 4]. Additionally, Fernandez et al. [5] highlighted the psychological impact of floods and suggested that floods can cause post-traumatic stress disorder, depression and anxiety, and long-term mental health implications for flood victims. It is the rare and extreme floods which have the most far-reaching consequences for both humans and the environment. Between 1998 and 2022 multiple extreme floods occurred in NSW and Queensland causing between \$900 million and \$3.4 billion losses. As well as causing fatalities, economic impacts continued for years afterwards.
Machine learning methods have been increasingly used to tackle time series prediction problems for a wide range of domains [6, 7, 8, 9] and are extensively used in hydrology [10, 11]. Machine learning methods can forecast streamflow at various lead-times, and show promise for hydrological forecasts of streamflow and floods [12, 13, 14, 15, 16, 17]. Hence, the literature motivates the use of machine learning models for extreme climate events such as floods in Australia.
A major challenge in hydrology is the ’regional modelling problem’, which is about the usage of either a model or a set of models to provide spatially continuous hydrological simulations across large areas (e.g., regional, continental, global). Regional models must learn and encode the similarities and differences of various catchment characteristics (such as topology, geometry, climate, ecology, etc.) and use them to meaningfully predict hydrological behaviour. Often this is done using ancillary data (e.g. soil maps, remote sensing, digital elevation maps, etc.) to map the similarity between catchments [18]. Kratzert et al. [19] showed how an long short-term memory (LSTM) [20, 21] model could be used to develop a single (regional) hydrologic model for hundreds of sites, concatenating the static data to the time-series data as inputs. LSTM models have outperformed traditional hydrological models for streamflow prediction in ungauged basins [22]. The compression of the static information used in hydrologic modeling can improve the integration of the static information with the time series data [23]. Ensemble learning methods use multiple machine learning models and combine their predictions for improved performance. In this case, the ensemble models complement the LSTM’s temporal information with the static features. Leontjeva et al. [24] showed that an LSTM model can extract information from temporal sequences, which can then be combined with ensemble methods like Extreme Value Machines and Random Forests to utilize static data. Miebs et al. [25] demonstrated that an ensemble with altered architectures of recurrent neural networks (RNNs) concatenated with dense networks could learn the temporal data with static supplementary information. Therefore, an ensemble machine learning model approach is promising, giving the opportunity for a regional approach to outperform a single-site approach calibrated per basin, which often suffers from inherent data shortage.
Forecasting time series with extremes is very difficult as the extreme values are often treated as outliers to improve the generalisation of the model. Most research on extreme values focuses on using extreme value theory (EVT) [26] as applications in anomaly-detection (detecting that an extreme event is present), rather than being able to predict the magnitude of the extreme event. An example is the work by Galib et al. [27] who used neural networks to predict the parameters of an extreme value distribution which consequently could predict the highest value within a specific time window of the time-series data.
Quantile regression models the conditional median (or other quantiles) of the response variable, as compared to linear regression, which estimates the conditional mean. Similarly, a quantile regression model training procedure minimizes the mean absolute error to estimate conditional quantiles of the response variable [28]. The quantile regression model is more robust to outliers in the response measurements and hence applicable to forecasting extremes [29, 30]. Quantile regression has been combined with machine learning models for several applications. Taylor et al. [31] introduced the quantile regression neural network, which offered a practical and flexible implementation compared to past work that involved linear or simple parametric non-linear models. Wang et al. [32] combined quantile regression with an LSTM model for load forecasting, and Pasche et al. [33] combined extreme value theory with quantile regression for a novel model which was applied to flood risk forecasting. Quantile regression found multiple applications in hydrology, such as in downscaling precipitation [34] and climate characteristics [35]. In relation to flood forecasting, Weerts et al. [36] applied quantile regression to assess the predictive uncertainty of rainfall-runoff and hydraulic forecasts. Cai and Reeve [30] addressed extreme and non-extreme value prediction via parametric quantile function models, showcasing differences among statistical models commonly used in extreme value prediction.
The domain of extreme flood prediction, particularly through the integration of deep learning with quantile regression, presents a unique opportunity to enhance the accuracy of flood forecasts. Despite significant advancements in forecasting streamflow and rainfall-runoff modelling, there is a critical gap in relation to predicting extreme flood events. This gap could be addressed by combining deep learning with quantile regression, which so far has not been used widely. This will specifically allow for assessing predictive uncertainty in hydrological forecasts. This study introduces an ensemble-based deep learning framework, which uses a switching mechanism inspired by extreme value theory for deep learning models, employing a multivariate and multi-step time-series prediction approach to forecast streamflow for multiple days ahead in major Australian catchments. We compare the performance of various deep learning models, including LSTM, bi-directional-LSTM, and CNNs, across a large number of basins. Our framework also features an ensemble model that integrates quantile regression with deep learning, enriching the time series with static information to support regional modelling. We provide open-source code along with data so that the methodology can be extended to other catchments in other regions.
The rest of the paper is organised as follows. In Section 2, we provide an overview of related topics, and in Section 3 we present our methodology. In Section 4, we provide the results, followed by a discussion in Section 5, and a conclusion in Section 6.
2 Background and Related Work
2.1 Hydrology
Surface water flooding is a result of accumulated rainfall covering a land surface since it cannot drain to natural or man-made drainage systems or watercourses due to them being at full capacity and overflowing [37]. In short duration or intense rainfall, flash flooding can arise from rapidly developing convective rainfall systems, which can be particularly dangerous in urban areas [38]. The prediction of flash floods for small to medium catchments remains a great challenge, as flash floods aren’t always caused by known meteorological phenomena [39]. Some factors that cause flash floods may include how quickly the storm is moving, the porosity of the soil, the steepness of the terrain, etc. Climate change is a big driver of extreme weather events that cause floods [40], and in particular rainfall events have been increasing in intensity [41].
The most common hydrological forecasting models are combined hydrological/hydrodynamic and meteorological models [42]. Hydrological models mostly simulate rainfall-runoff relationships, while hydraulic models deal with channelflow and flow over surfaces. Meteorological rainfall forecast models are needed, as ultimately rainfall drives floods, but these need to quantify uncertainty [43]. The problem is that there are numerous sources of uncertainty, which limits forecasting ability [44]. The general approach to manage uncertainty has been to use ensemble forecasts. For example, Emerton et al. [45] used ensemble streamflow forecasting and featured a number of forecast scenarios to explicitly characterize forecast uncertainties. At a much larger scale, Alfieri et al. [46] presented a global flood awareness system (GloFAS) for detecting the probability of streamflow given warning thresholds by using ensemble models. However, Cloke et al. [47] outlined that ensemble models can be limited by computational speed and complexity. Ultimately, hydrological models are constrained by the amount of information available, such as the input of meteorological variables (e.g. flow, water level, rainfall) and the need to parameterise geological, topographic and soil variables [48].
As mentioned, most hydrological models are calibrated for a single catchment and the challenge is to extrapolate to different catchment across a region. In regionalisation approaches, model-dependent (parametric) and model-independent (non-parametric) approaches are the two primary strategies [18].
The model-dependent approaches are those that assume a pre-defined hydrological model across all catchments, wherein the model parameters are to be estimated from the data. For example, different approaches can be used for regionalisation, either focusing on spatial proximity of or on hydrological similarity [49], and spatial proximity gave the best performance across catchments in France. However all approaches were outperformed by models calibrated to single catchments [49]. In a recent uncertainty focused approach, Prieto et al. [50] aimed at identifying not only model parameters, but different model structures in a regionalisation approach.
Conversely, model-independent methods, non-parametric in nature, require no prior assumptions and are driven entirely by the data [22]. It identifies the mapping between ancillary data, catchment attributes and other inputs to predict streamflow. The most common of these are machine learning models such as neural networks, being most effective with large amounts of data, and allowing knowledge to be transferred between sites [51]. Several papers [52, 19, 53] have shown that LSTM approaches in particular can outperform a numerical conceptual hydrological models,even if they are calibrated individually for each catchment in a gauged setting [52].
2.2 Deep learning and ensemble models
The success of recursive neural networks (RNN) has been mostly demonstrated in language models [54], but it has also been prominent for modelling temporal sequences [55] and multistep ahead prediction [56]. Some of the major categories of time series forecasting (prediction) include one-step, multi-step and multi-variate time series prediction. The one-step ahead time series prediction refers to models that have a prediction horizon (step) of a single time unit, whereas multi-step predictions refers to models that have a prediction horizon of multiple time units. The challenge is in developing models that maintain good prediction accuracy as the prediction time step increases, since the chaotic nature and noise of time-series data makes this difficult [57, 56].
RNN models cannot learn long-term dependency, as these do not have "memory" to capture autocorrelation present in long streams of the data [58]. This limitation was addressed by the long short-term memory (LSTM) network [20] through the inclusion of internal memory cells that control the flow of information in the model. This has improved natural language processing (NLP) [59] and language translation [60] which involve longer-term dependencies, since they have data with long sequences, i.e a paragraph in a text or the data can be the entire chapter which needs to be translated.
Bi-directional RNN and LSTM (Bi-RNN, Bi-LSTM) further address the shortcomings of earlier counterparts, where only direct previous context states are used for determining future states. The bi-directional architecture consists of placing two independent RNN/LSTM layers together to allow information to be processed in both directions of the network, at every time step [61]. Accordingly, information is transferred from past to future and from future to past[62].
Convolutional neural networks (CNN) were designed to learn spatial hierarchies of features through multiple building blocks of convolution, pooling and fully connected layers. Convolution layer(s) in a CNN were originally designed for image classification purposes [63], but can also be used for temporal sequences, including time series prediction [64, 65, 56]. For example CNN has been used for precipitation downscaling [66].
Suppose $Y$ is the real-valued random value for the response, then $F_{Y}(y)=P(Y\leq y)$ represents the cumulative distribution function. The $\tau^{t}h$ quantile for $\tau\in(0,1)$ may be given by:
$q_{Y}(t)=F_{Y}^{-1}(\tau)=inf\{y:F_{Y}(y)\geq\tau\}$ (1)
Rather than using the squared error-based function, we define the tilted absolute value function as $\rho_{\tau}(m)$ for a particular quantile $\tau$.
$q_{Y}(\tau)=arg_{u}minE(\rho_{\tau}(Y-u))$ (2)
As highlighted earlier in the introduction, deep learning and particular LTSM models can capture memory and therefore autocorrelation, as well as linear and non-linear behaviour in hydrological systems. LSTM models can learn long-term dependencies for temporal sequences and exhibit better performance for flood forecasting [67] when compared to simple neural networks. Although the CNN[64, 65] was designed for computer vision tasks, they also have shown remarkable results in time-series prediction. In particular, Chandra et al. [56] showed that CNN is competitive with LSTM for multi-step ahead time series prediction. Furthermore, Hussain et al. [68] implemented a one-dimensional (1D) CNN for one-step-ahead streamflow forecasting for three-time horizons (daily, weekly and monthly) in Gilgit River of Pakistan. Khosravi et al. [69] used CNN with meta-heuristic optimisation and reported superior predictive performance in daily stream-flow forecasting compared to related models.
Similarly highlighted in the introduction, ensemble learning models combine multiple machine learning methods as base models to produce an optimal predictive model [70, 71]. Some of the prominent ensemble learning models are bagging and boosting via implementations such as random forests and adaboost [72]. In hydrology, ensemble models have also been applied extensively [11], where the main question has focused on how to construct the ensemble. Should this by averaging methods [73], bootstrapping [74] or stacking [75, 76]. Initially, Cannon et al. [77] investigated how climate change impacted streamflow using ensemble neural networks and multiple linear regression methods. Further work by Diks et al. [73] compared several ensemble methods such as equal weights averaging, and indicated that the Granger-Ramanathan averaging method most improved forecast accuracy for environmental systems. In contrast, Sharma et al. [74] demonstrated that a hybrid wavelet bootstrapped neural network provided better predictions than conventional machine learning models in forecasting daily river discharge. Tyralis et al. [75] took a different approach, to forecast short-term daily streamflow, by stacking ten different machine learning models. Furthermore, Li et al. [76] showed an example, which integrated stacked models such as support vector machines, elastic-net regression and extreme gradient boosting (XGBoost) model to forecast midterm stream-flow. However, bagging has been the predominant ensemble model until 2010 in hydrology, but more recently, boosting and stacking has dominated due to improved performance [11].
3 Data and Methodology
3.1 Framework
Based on the reviewed literature, the goal of this study is:
  1. 1.
    Assess the ability of selected deep learning models to accurately reproduce streamflow estimates for the given catchments.
  2. 2.
    Examine several methods of introducing static catchment attributes as additional inputs, such as the Batch-Static and Stacked Ensemble Architecture.
  3. 3.
    Compare selected deep learning model and data configuration to evaluate model strengths and failures when attempting to predict extreme rainfall events such as floods.
  4. 4.
    Address the failure of conventional models by introducing a novel conditional ensemble deep learning model that employs features of quantile regression.
Figure 2 presents the framework, which features three key stages. Using our framework, we examine deep learning models with catchment data strategies as given below:
  1. 1.
    The Individual strategy features a separate model for each base station, as shown in Figure 4.
  2. 2.
    The Batch-Indicator strategy includes base stations from the entire region, where data from all the basins are concatenated for a single model as shown in Figure 5.
  3. 3.
    Batch-Static strategy features data from the entire region with static input, where each input vector features static variables as shown in Figure 6.
  4. 4.
    The Ensemble catchment strategy (Figure 7) combines forecasts from two models: one capturing temporal sequences and the other static features. These forecasts are concatenated and processed through a regression layer to generate final predictions. We investigate if the overall prediction accuracy improves by leveraging the strengths of both temporal and static feature models, enhancing the model’s ability to make more accurate and robust predictions.
In the initial stage, we assess the performance of four distinct model architectures: i.e. an Individual architecture, a Batch Indicator, a Batch Static, and an Ensemble model. Then we advance to the second stage, where we evaluate the performance of four deep learning models within the chosen architecture, based on the summary provided in 2.2: LSTM, CNN, BD-LSTM, and ED-LSTM. In the third stage, we utilise the conditional ensemble model that incorporates features from quantile regression. We compare it against the previously identified best-performing models to derive valuable insights about predictive capabilities and reliability.
3.2 CAMELS dataset
We make use of the catchment attributes and meteorological (CAMELS) dataset [78] to demonstrate the use of deep learning models in flood predictions. CAMELS is a set of data containing hydrometeorological time series and landscape attributes for 222 catchments in Australia, Figure 1 presents an overview of the approximate locations.
Figure 1: Location of the Camels catchments across Australia across the different states.
Table 1: Summary of available data in CAMELS dataset
Type Category/Description Units
Time Series Streamflow ml $d^{-1}$
Time Series Precipitation mm $d^{-1}$
Time Series Actual and potential evapotranspiration mm $d^{-1}$
Time Series Maximum Temperature °C
Time Series Minimum Temperature °C
Summary Mean streamflow mm $d^{-1}$
Summary Sensitivity of streamflow to rainfall -
Summary Ratio of streamflow to precipitation -
Summary Frequency of high flow days d y-1
Summary Average duration of high flow days
Summary Frequency of low flow days d y-1
Summary Frequency of days with zero flow d y-1
The CAMELS dataset is suitable for “large-sample hydrology” [79, 80] and for evaluation of deep learning models [81]. The CAMELS time series data consists of streamflow and 18 climatic variables at daily timesteps. The start of the time series is the 1st of January, 1911; however, since more stations were added over time, the starting date of each time series per station is different. For example, for Station A, streamflow data may range from 1 January 1950 to 1 January 2020 , but for Station B streamflow data may range from 1 January 1980 to 1 January 2020. As a result, if the analysis is across stations , we must take the dates that are all in common.
The landscape attribute gives data and descriptions of attributes related to geology, soil, topography, land cover, anthropogenic influence and hydroclimatology for all 222 catchments. This is static information, so each catchment will only have one value for each attribute. The metadata is also available, which describes the latitude, longitude, boundaries of each catchment. We provide a summary of the different data types in Table 1.
The dataset contains a variety of data that may not be relevant to the problem of flood prediction, so we select only the relevant time series and static variables. In all our models, the input time series includes i.) daily cumulative precipitation, ii.) daily minimum air temperature, iii.) daily maximum air temperature, iv.) the time of year as sine/cosine functions.
Furthermore, we selected certain catchment characteristics for the models that require static data. These include i.) mean streamflow, ii.) sensitivity of streamflow to rainfall. iii.) ratio of streamflow to precipitation. iv.) frequency of high flow days. v.) average duration of high flow. vi.) frequency of low flow days. vii.) frequency of days with zero flow. These catchment attributes include climatic and vegetation indices, as well as soil and topographical properties.
3.3 Data Reconstruction
In order to use deep learning models, the original time-series data (which is common in climate and hydrological science) needs to be transformed (embedded) to be used for either single-step or multistep-ahead prediction. For this, the data needs to be processed using a sliding window approach, and the size of the window is crucial for this step. Taken’s theorem demonstrates that any smooth coordinate transformation does preserve the important properties of the original time-series data [82] with an appropriate embedding dimension. Thus, if the original univariate time-series data is of the form:
$Y=[x(1),x(2),...,x(t),...,x(N)]$ (3)
where, we observe $x(t)$ at time $t$. We can create an  textitembedded phase space via sliding windows:
$Y(t)=[x(t),x(t-T),x(t-2T),...,x(t-(D-1)T)]$ (4)
where, $T$ is the time delay, $D$ is the embedding dimension (window-size); this is valid for all values of $t$ from $t=0,1,2,...,N-DT-1$ and $N$ is the length of the data. Assuming the original attractor is of dimension $d$, then $D=2d+1$ would be sufficient to capture the important characteristics of the original data in the embedded phase space [82]. We will use the embedded phase space for training the deep learning models where $D$ will be used to define the input size (neurons) of the respective deep learning models. Essentially, a univariate time series vector gets transformed into a matrix that is used for training and testing the model. Rather than using the entire univariate time series, the embedded phase space enables a sliding window approach for training the respective deep learning model, where the input is defined by the size of the sliding window. In the case of a multivariate time series, the same approach can be used, as shown in Figure 3.
We need to transform the two-dimensional dataset of features and temporal sequences (daily time interval) into a three-dimensional format through a sliding window technique. A window of size of $N$ slides over the 2-dimensional dataset given by $6\times T$; where 6 is the number of input variables, and $T$ is the length of the dataset (number of days for available data). This creates a 3D dataset of $M\times N\times 6$; where $M=T-N+1$, i.e. the number of times the window can slide over the dataset. As a result, we create a feature input dataset $X_{i}=[x_{1},x_{2},...,x_{5}][s_{1}],\;i=1...N$ and target (prediction) $Y_{j}=[y_{1},y_{2},...,y_{N_{o}ut}],\;j=1...N-out$, as shown in Figure 3.
Figure 2: Framework for comparing the different models for the different architectures.
Figure 3: Embedding: transformation of a two-dimensional temporal dataset of features (daily time interval) into a three-dimensional format through a sliding window technique.
Figure 4: Individual model per station. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction). …
Figure 5: Batch model. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction).
Figure 6: Batch model with static data. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. The arrangement of [s1] is for models requiring static features as a one-time input. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction).
Figure 7: Stacked Ensemble strategy combining Temporal and Static models using Linear Regression.
3.4 Quantile-LSTM: conditional ensemble deep learning model
Our conditional ensemble deep learning model (Quantile-LSTM) capitalises on quantile regression using a 1D-LSTM model (Figure 8). We will use the same data as for the deep learning models for prediction of the flow-duration exceedance probability ($\alpha$), which is the cumulative frequency curve that shows the probability of streamflow exceeding a given threshold for the data period [83]. In essence, this initial model predicts the probability that a flood of a certain magnitude occurs in the prediction window.
Quantile-LSTM is based on the quantiles of a distribution, and will offer a comprehensive understanding of prediction uncertainty. We strategically use three separate Quantile-LSTM models to optimize the performance for different flood probability scenarios. Depending on the predicted probability, the LSTM models provide three separate n-step-ahead predictions; i.e. we train a Quantile-LSTM at a quantile level ($q=0.95$), when the predicted flood exceedance probability is above a threshold ($\alpha>0.95$), ensuring accurate forecasting in high-probability flood events. Similarly, if $\alpha$ falls between 0.7 and 0.95, we use a Quantile-LSTM with $q=0.70$. In contrast, when $\alpha<0.70$, we train a standard LSTM model since this region does not feature extreme values in data. This tripartite approach allows the model to adapt to varying flood risk levels, capitalising on the flexibility and specificity of the Quantile-LSTM in capturing different aspects of the flow-duration curve. Figure 8 presents the details of the switching mechanism for the conditional ensemble model which decides which Quantile-LSTM model to be used. We will evaluate this approach against the observed data and the best-performing deep learning models.
Figure 8: Conditional ensemble model utilising Quantile-LSTM.
3.5 Experiment Setup
In this section, we provide technical details for the baseline deep learning models (LSTM, BD-LSTM, ED-LSTM, and CNN) and our proposed models that include Stacked Ensemble and Quantile-LSTM. We determined model hyperparameters such as hidden neurons, learning rate and epochs from trial experiments, as shown in Table 2. We note that all the tested models implement a multi-step time series prediction with 5 inputs and 5 outputs (time steps). We train the model with a maximum time of 150 epochs using an Adam optimiser [84] with default hyperparameters, such as the initial learning rate.
In order to develop regional-based models, we separate the CAMELS data based on the state. All time-series data are scaled using a min-max-scaler, i.e. scaled in the range [0,1]. In each basin, there are about 10,000 data points representing daily data spanning from 1990 to 2015. We use 60% for training and 40% for testing. We used the first 60 % of the data from 1990 as training data and the remaining for testing. We note that although there may be a climate bias in taking a block of data spanning more than a decade, this approach ensures reproducibility in results when compared to shuffling the data.
It is necessary to use a variety of performance metrics for model benchmarking, as no one evaluation metric can fully capture the consistency, reliability, accuracy, and precision of a stream flow model. The Nash–Sutcliffe-Efficiency (NSE) is the hydrology version of coefficient of determination, similar to $R^{2}$, which can be less than zero. Silva et al. [85] have suggested the Squared Error Relevance (SER) - (SER-$\alpha$%) for extreme events. The SER calculates the Root Mean Squared Error (RMSE) of test windows that contain the $\alpha$ percentile of the data, measuring the ability of a model to predict rare data points. For example, SER-1% would find the RMSE of windows that contain the 1% most extreme stream flows for a given basin. In summary, we use the following metrics to evaluate prediction performance accuracy over multiple-steps (prediction horizon).
  1. 1.
    The RMSE is given by $\sqrt{\sum_{i=1}^{N}(\hat{y_{i}}-y_{i})^{2}}$ between predicted $\hat{y_{i}}$ and actual $y_{i}$ values for $N$ data points.
  2. 2.
    The SER is given by $i\%$ RMSE $\sqrt{\sum_{i\in D^{i}}(\hat{y_{i}}-y_{i})^{2}}$ of the windows that contain stream flows above the ith percentile in the test dataset, $i\in(1,2,5,10,25,50,75)$.
  3. 3.
    The NSE is given by $1-\frac{\sum_{i=1}^{N}(Q^{t}_{o}-Q^{t}_{m})^{2}}{\sum_{i=1}^{N}(Q^{t}_{o}-\bar{Q^{t}_{o}})^{2}}$ and ranges from ($\infty$, 1).
Table 2: Model Configuration
Models Hidden Neurons Comments
LSTM 20 Cells Single Layered
BD-LSTM 20 Cells Single Layered
ED-LSTM 20 Cells Single Layered
CNN 64 Cells Single Layered
Quantile-LSTMq=0.5 20 cells Conditional Ensemble Model
Quantile-LSTMq=0.7 20 cells Conditional ensemble Model
LSTM 20 cells Conditional Ensemble Model
Table 3: Comparison of catchment strategies using LSTM model for South Australia
Architecture SER 1% SER 2% SER 5% SER 10% SER 25% SER 50% SER 75% RMSE
Individual 1.8321 1.1904 0.6606 0.4165 0.2226 0.1337 0.1064 0.0963
±0.0412 ±0.0221 ±0.0143 ±0.0117 ±0.0082 ±0.0072 ±0.0069 ±0.0068
Batch-Indicator 2.2775 1.6474 1.0696 0.6884 0.3669 0.2332 0.1915 0.1789
±0.1107 ±0.0884 ±0.1174 ±0.0670 ±0.0343 ±0.0229 ±0.0203 ±0.0200
Batch-Static 1.9712 1.2450 0.6755 0.4853 0.2312 0.1461 0.1127 0.1040
±0.2421 ±0.1556 ±0.0805 ±0.0459 ±0.0227 ±0.0164 ±0.0134 ±0.0121
Stacked Ensemble 4.1612 2.6328 1.3091 0.7229 0.3363 0.1975 0.1627 0.1499
±0.0445 ±0.0322 ±0.0181 ±0.0109 ±0.0050 ±0.0024 ±0.0017 ±0.0014
Table 4: Architecture wise- Model Comparisons to track best performing model (South Australia)
Model SER 1% SER 2% SER 5% SER 10% SER 25% SER 50% SER 75% RMSE
Individual LSTM 1.8321 1.1904 0.6606 0.4165 0.2226 0.1337 0.1064 0.0963
±0.0412 ±0.0221 ±0.0143 ±0.0117 ±0.0082 ±0.0072 ±0.0069 ±0.0068
Individual BD-LSTM 1.7041 1.1264 0.6426 0.4029 0.2178 0.1384 0.1143 0.1053
±0.0529 ±0.0383 ±0.0309 ±0.0222 ±0.0123 ±0.0092 ±0.0089 ±0.0089
Individual CNN 1.7349 1.1204 0.6199 0.3776 0.1964 0.1273 0.1056 0.0976
±0.0564 ±0.0288 ±0.0158 ±0.0085 ±0.0035 ±0.0018 ±0.0016 ±0.0016
Individual ED-LSTM 1.7468 1.1528 0.6508 0.4071 0.2188 0.1390 0.1144 0.1055
±0.0251 ±0.0183 ±0.0199 ±0.0168 ±0.0121 ±0.0100 ±0.0094 ±0.0092
Quantile-LSTM 1.6509 1.0572 0.5826 0.3604 0.1963 0.1220 0.1006 0.0929
±0.0743 ±0.0418 ±0.0181 ±0.0093 ±0.0040 ±0.0021 ±0.0014 ±0.0011
Table 5: Quantile-LSTM results for major states in Australia
State Number of Stations SER 1% SER 2% SER 5% SER 10% SER 25% SER 50% SER 75% RMSE
SA 9 1.6509 1.0572 0.5826 0.3604 0.1963 0.1220 0.1006 0.0929
±0.0743 ±0.0418 ±0.0181 ±0.0093 ±0.0040 ±0.0021 ±0.0014 ±0.0011
NSW 33 1.4779 0.9425 0.4808 0.2770 0.1285 0.0695 0.0490 0.0401
±0.2421 ±0.1556 ±0.0805 ±0.0459 ±0.0227 ±0.0164 ±0.0134 ±0.0121
QLD 61 1.4556 0.9937 0.5355 0.3117 0.1396 0.0727 0.0506 0.0425
±0.2408 ±0.1519 ±0.0961 ±0.0678 ±0.0476 ±0.0263 ±0.0214 ±0.0122
VIC 72 1.0603 0.6685 0.3685 0.2241 0.1165 0.0693 0.0528 0.0451
±0.2817 ±0.1823 ±0.1041 ±0.0704 ±0.0502 ±0.0286 ±0.0230 ±0.0134
WA 16 2.3148 1.7796 1.1950 0.8078 0.4384 0.2735 0.2132 0.1891
±0.1550 ±0.0827 ±0.0402 ±0.0331 ±0.0193 ±0.0133 ±0.0118 ±0.0113
NT 16 1.7982 1.2585 0.7505 0.4838 0.2431 0.1509 0.1241 0.1104
±0.0251 ±0.0183 ±0.0199 ±0.0168 ±0.0121 ±0.0100 ±0.0094 ±0.0092
TAS 12 1.4647 1.0495 0.6412 0.4272 0.2419 0.1526 0.1129 0.0876
±0.0791 ±0.0591 ±0.0379 ±0.0243 ±0.0145 ±0.0119 ±0.0092 ±0.0081
ACT 3 1.7321 1.0715 0.4775 0.2528 0.1147 0.0607 0.0413 0.0317
±0.0564 ±0.0288 ±0.0158 ±0.0085 ±0.0035 ±0.0018 ±0.0016 ±0.0016
4 Results
As outlined in section LABEL:label:Framework, we evaluate conventional deep learning models using performance metrics which take into account the catchments, prediction horizon, and window length of the input time series. We evaluate four distinct catchment modelling strategies using a conventional LSTM model to identify the optimal catchment modelling strategy using a subset of the data from South Australia as a case study. We compare the Individual, Batch-Indicator, Batch-Static and Ensemble catchment modelling strategies. The analysis considers metrics such as SER(with percentile cuts at 1, 2, 5, 10, 50, and 75) and RMSE mean across five prediction horizons. This enables a thorough evaluation the influence of extreme flood conditions while making comparisons across different models and architectures. Table 3 presents the results, highlighting that the Individual catchment strategy emerged as the most effective across various metrics (SER percentiles). This strategy has the lowest (best) RMSE, which indicates that separate models for individual gauging stations gives the best performance. This is because each catchment has unique characteristics that could be based on weather patterns, climate, and topography. In Table 3, the Stacked Ensemble model gives poor performance, which is not even able to beat the individual LSTM models. This could be due to problems in combining the data from the Temporal with the Static model.
The next step is to compare different deep learning models for catchments in South Australia, including the LSTM, BD-LSTM, CNN, ED-LSTM, and Quantile-LSTM. We chose South Australia dataset for comprehensive evaluation since it has a lower number of gauging stations and data when compared to large states such as New South Wales. As shown in Table 4, we find that the Quantile-LSTM provides the best prediction accuracy (RMSE of 0.0929) and SER percentiles. This is close to the LSTM model (RMSE of 0.0963), and CNN (RMSE of 0.0976), followed by BD-LSTM and ED-LSTM which give the lowest performance, respectively.
Based on the analysis in South Australia, the Quantile-LSTM provides the best results and thus we proceed to employ it across all states in Australia. Table 5 presents the results, highlighting that in some states, the prediction accuracy (RMSE) is better than others, which is correlated to the SER percentiles. The difference in the results across the states could be due to different levels of extreme values in the data, and also the size of the data and the presence of noise. The Northern Territory and Western Australia have the worst model accuracy (RMSE) while New South Wales, Australian Capital Territory, and Queensland, and Victoria have some of the best prediction accuracies (Table 5). Notably, New South Wales and Queensland feature the highest number of catchments (gauging stations) and are states that are prone to extreme floods.
Finally, in Figures 12 to 16, we provide a visualisation of actual and predicted values using Quantile-LSTM for selected gauging stations (using Basin ID) for the different states for a period of more than two years. In most of the figures, we notice that short duration extreme rainfall events that contribute to high flood probabilities are predicted well by the Quantile-LSTM model. However, in most cases (eg. Figure 9 2010-07 - 2010-10 and Figure 12 2012-04 - 2012-07), the Quantile-LSTM model provides a lower flood probability when the actual flood probability is very large. This is noticeable in most of the selected Basin IDs for the respective states. This bias in prediction can potentially be corrected with a weighting mechanism, data augmentation or regularisation mechanism in the deep learning models.
Figure 9: South Australia Basin ID: A5030502
Figure 10: New South Wales Basin ID: 401009
Figure 11: Queensland Basin ID: 919309A
Figure 12: Victoria Basin ID: 401203
Figure 13: Australian Capital Territory Basin ID: 410730
Figure 14: Western Australia Basin ID: 604053
Figure 15: Northern Territory Basin ID: G8200045
Figure 16: Tasmania Basin ID: 318076
5 Discussion
In the first set of experiments, we compared the performance of the modelling strategies at the catchment scale using a conventional LSTM model for five step (day)-ahead prediction of streamflow (flood probability) using climate data. We found that the Individual catchment modelling strategy, where a separate model is used for each gauging station data, provided the best results (Table 3). This is because each catchment has unique characteristics that could be based on weather patterns, climate, and topography. In Table 3, the Stacked Ensemble model gives poor performance, which is not even able to beat the individual LSTM models. This could be due to problems in combining the data from the Temporal with the Static model. Generally, it is difficult for a single machine learning model to represent a wide range of data, even if they may be from the same domain, such as in a regionalization context. Regionalization would be a challenge for machine learning in general, as a single model cannot be used (fitted to) for several tasks, and in this case, the local catchments can be seen as different tasks as they can have differences in the data distribution dependent on climate/weather/geography. Approaches such as multitask and transfer learning [86, 87] could be considered, where some underlying model knowledge can be re-used and applied to related datasets, i.e. catchments in this case.
In the case of deep learning models using the Individual catchment strategy (Table 4), we observe that the Quantile-LSTM model provides the best results. However, Conventional LSTM and CNN models provide accuracy close to Quantile-LSTM model, which specifically includes a mechanism for quantifying uncertainty. Although designed for image processing purposes, the CNN reproduces streamflow better than the LSTM models that were designed for temporal sequences (it captures windows well that have little streamflow, as well as windows that contain extreme streamflows). CNNs may also have advantages for capturing spatial regions (regionalisation) due to their capabilities in image processing, but LSTM models were the best due to their capacity to simulate temporal data. The performance of these types of models can however be comparable as was shown in previous studies, especially for multi-step ahead prediction [56]. Furthermore, the Individual catchment strategy does not incorporate the spatial catchment data into the deep learning models, although multivariate data of the local catchment was used. Hence, this could be another reason why LSTM models outperformed the CNN. ED-LSTM and BD-LSTM, in this case, did not provide the a good performance, as was highlighted in past studies for univariate multi-step ahead prediction [56]. This could be due to the multivariate nature of the data that features extremes.
The results, reinforce the efficacy of the Quantile-LSTM as a robust and adaptable choice for hydrological modelling, specifically for South Australia. This performance highlights the Quantile-LSTM model resilience and effectiveness in capturing both extreme and regular flood occurrences, thereby reinforcing its suitability for capturing long-term hydrological patterns. However, as indicated, the Quantile-LSTM model can be biased in the extremes. This bias in prediction can potentially be corrected with a weighting mechanism, data augmentation or regularisation mechanism in the deep learning models.
The prediction of the future steps essentially gives an outlook - up to five days ahead, which would be useful for extreme events. However, the performance accuracy deteriorates as the prediction horizon increases in multi-step ahead problems (RMSE, SER). The results deteriorate as the gap in the values to predict increases with increasing prediction horizon.
In our initial experiments (Table 3), we found the Individual catchment strategy to have the best performance. Interestingly, the Batch-Static strategy uses the same static features (7 summary features, stated in Table 1 ) as the Ensemble strategy which performs much worse (RMSE and SER percentiles). It seems that the Stacked Ensemble strategy cannot meaningfully extract and combine the static and temporal information from the respective models (Table 3). Comparatively, The Batch-Indicator Model is essentially the Batch Model without the inclusion of static features. It operates on the same time-series features as the Individual Architecture, but with a collective approach, running a single model for all stations in the state rather than individual models for each station. This consolidation aims to optimize computational resources.. The Stacked Ensemble strategy first generated activations from the temporal data using one model, and then generated activations from the static data in another model to more accurately capture temporal and static information in an ensemble. The Individual strategy features a specific fit to one catchment at a time and is unaffected by information from other catchments. For example, catchments might vary due to variations in weather patterns, climate, and topography. Overall, these findings are consistent with the work of Leontjeva et al. [24] who similarly compared a variety of strategies i) standalone model using univariate temporal data ii) standalone model using bimodal data (temporal and static) iii) ensemble model using bimodal data. In that study the standalone model using bimodal data fails to capture information from both sources, while the performance of the ensemble model depended on the distribution of the data.
Despite the success of evaluating several machine learning approaches for predicting flood extremes, this study has some obvious limitations. Firstly, the models do not naturally project model uncertainties into the predictions, which is vital for natural disasters caused by extreme events. This can be addressed by Bayesian deep learning [88, 89]. Secondly, we have not accounted for topography information in the data to develop a robust spatiotemporal model.
In the case of larger catchments, the steepness and variation in topography can influence rainfall patterns, and particularly the within catchment distribution of the rainfall. In addition, topography and catchment shape can influence the timing between rainfall and flood extremes. Implicitly, the LSTM models should be able to capture these variations for the individual catchment strategy based on the data, but it is possible that adding additional information will assist the model in identifying the patterns. We have also did not incorporate the changes in climate conditions due to El Nino climate effect, which is a longer scale pattern in the Australian rainfall [90, 91]. Again, the LSTM should be able to discover this from the data, given a long enough time series. However, adding this as a variable might assist the model identifying the variations in the data. Finally, the overall data set does not have the same length and missing data across the individual stations. The data is also sparse in space given the locations of the gauging stations (catchments), and this means if we need data for specific regions where no gauging stations are available, generative machine learning models [92], and general circulation models [93] could be used.
6 Conclusions
In this comprehensive study, we undertook a critical evaluation of existing deep learning models and introduced innovative methodologies for the accurate prediction of streamflow extremes across numerous catchments. Leveraging a multi-step-ahead forecasting strategy, our investigation revealed that a tailored approach, where each catchment is paired with a bespoke model dedicated to its distinct gauging station, significantly elevates the accuracy of extreme value forecasts. It’s important to highlight that this precision in modeling ensures that local variability and unique hydrological characteristics are captured effectively, providing a more nuanced understanding of each catchment’s behavior in the face of extreme weather events.
Moving to the forefront of our findings, we unveiled the potential of our conditional ensemble model, named Quantile-LSTM, which integrates quantile regression to surpass all baseline machine learning models in terms of predictive error margins for streamflows. This model emerges as a superior tool not only due to its performance but also because it represents a leap towards a more reliable prediction by embracing the inherent uncertainties in hydrological forecasting. Looking ahead, we recognize that the path to impeccable forecasts is iterative and progressive. Our research underscores the vast room for improvement that can be addressed by extending our dataset and refining our model, particularly in the area of uncertainty quantification. The aspiration is to reach a point where our model can offer robust predictions with quantified confidence levels, transforming how stakeholders prepare for and respond to the challenges posed by climate-induced flood extremes.
Software and Data Availability
We provide meta-information for open source code and data for the framework via Github repository:
    Name: Streamflow-floods
    Developers: Siddharth Khedkar and Rohitash Chandra
    Contact email: siddharthkhedkar11@gmail.com and c.rohitash@gmail.com
    Compatible Operating System: Mac/Linux/Windows
    Developed and tested: Ubuntu 20.04 (Linux)
    Year published: 2024
Author contributions
S. Khedkar and J. Ng contributed to coding, writing and experiments. R. Vervoort contributed to conceptualisation, writing and supervision. R. Chandra contributed to conceptualisation, writing, code analysis, and supervision.